Consider a production function of the following form: F(Kt, Nt) = ln(φKt + θln(Nt)) where Kt is physical capital, Nt is labor hours, and θ,φ are exogenous parameters. Does this production function satisfy basic economic theories? Explain/prove why or why not
Answer #1

The production fenation is F (kt, Nt) = in [ukt + ln (NE)] = 4+ MP = 2 = 1 kk + 8 in (Ne)] * * Y+ MP4 = . xe TWKŁ + oln (NU)

Here the marginal productivity of both inputs are positive implied both the inputs are positively related to the production function, => as any of these inputs increases the production increases. Now, both the marginal products are negative implied as more and more of additional inputs are added the output increases but at a decreasing rate, => “diminishing marginal productivity” holds. So, the given production function holds all the economic theories.